Circular No. 7747

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SS 433 T. Kotani, Goddard Space Flight Center; and S. Trushkin, Labora-

tory of Radio-Astronomical Observations, Special Astrophysical Observatory, write: “The RATAN-600 Radio Telescope (flux density limit 0.5 mJy; absolute brightness temperature limit 0.05 mK) in the North Caucasus, Russia, has observed a flare in the microquasar SS433 on Nov. 2. Flux densities reached 1.3 Jy at 2.3 GHz on Nov. 3.6 UT. The flare indicates that the source entered its high state, and successive flares exceeding 1 Jy with intervals of 8–23 days are expected. The probability of the occurrence of a second flare within 23 days is estimated to be 90 percent, based on the archival radio data. The source will be monitored in the radio band until Dec. 31, and PCA/RXTE monitoring observations will begin on Nov. 9.3. If SS433 experiences a flare during this period, it will be the first flare from the source observed in a multi-wavelength campaign. We strongly encourage spectroscopic and photometric monitoring observations.”

1RXS J232953.9+062814 M. Uemura, R. Ishioka, and T. Kato, Kyoto University; P. Schmeer,

Bischmisheim, Germany; H. Yamaoka, Kyushu University; D. Starkey, Auburn, IN; T. Vanmunster, Landen, Belgium; and J. Pietz, Erftstadt, Germany, on behalf of the VSNET collaboration team, write: “On Nov. 3.926 UT, P. Schmeer detected an outburst (at m_{v} = 12.5) of this object, classified as a dwarf nova by Jingyao et al. (1998, Ann. Shanghai Obs. 19, 235). Our CCD photometry (bandpass close to Cousins R) on Nov. 4.47– 6.17 revealed superhumps with amplitudes of 0.2–0.3 mag and a period of 0.046311(12) day, indicating that this object is an SU UMa-type dwarf nova. This short superhump period means that the orbital period of this object is certainly below the ‘period minimum’ of hydrogen-rich cataclysmic variables (∼ 1.3 hr). According to Jingyao et al., the quiescent spectrum was dominated by Balmer emission lines, which indicates that this object is not an AM CVn star but a hydrogen-rich system. Except for this object, we know only one source (V485 Cen) in this class (IAUC 6666). Comparing our CCD images with Digital Sky Survey images, this object shows noticeable proper motion (up to 0′′.1/yr). Its relative brightness and large proper motion indicate a small distance for this object, which is thus one of the most important systems to study the evolutionary scenario of cataclysmic variables. Follow-up observations are strongly encouraged.”

2001 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green